30. Solve: *+2_ 30 Show work. X-3
Answer #1

3 30 2

30 2-3 (г. +3)(z-3)

113 I(r+3)

\Rightarrow (x+2)(x+3)=30

25r 6- 300

\Rightarrow x^{2}+5x-24=0

\Rightarrow x^{2}+8x-3x-24=0

\Rightarrow x(x+8)-3(x+8)=0

(z+8)(2-3) = 0

Therefore r=-8, 3